# CREME: Robustness Enhancement of Code LLMs via Layer-Aware Model Editing

### v2

@@ 1. Introduction
In recent years, the rapid development of large language models (LLMs) has led to the emergence of powerful models, including ChatGPT (Achiam et al., 2023), LLaMA (Touvron et al., 2023), and DeepSeek (Liu et al., 2024). Trained on large-scale textual corpora, these models exhibit strong generalization capabilities and have achieved notable success across a wide range of software engineering tasks. Among these tasks, code generation has emerged as a key application in AI-assisted software engineering, attracting growing research attention (Chen et al., 2021; Li et al., 2023; Touvron et al., 2023; Lin et al., 2024; Chen et al., 2024a). Typically, practitioners employ LLMs by providing natural language descriptions and LLMs generate the source code, which automates programming tasks and accelerates development workflows.
@@
The natural language description in a prompt is crucial for conveying the requirements defined by users to LLMs. Prior studies (Du et al., 2023; Chen et al., 2021; Austin et al., 2021) have evaluated the code generation capabilities of LLMs using datasets consisting of human-verified prompts. However, in real-world scenarios, prompts submitted to LLMs often differ in wording, syntax, and formatting. They may also contain typographical errors or redundant expressions. Prior studies (Shen et al., 2023; Wang et al., 2023; Chen et al., 2024b; Shirafuji et al., 2023) have shown that LLMs are sensitive to such minor variations, even a slight change may lead to a completely different result. Not all users of LLMs are skilled prompt engineers capable of making precise, error-free prompts. Therefore, it is essential to ensure output stability when semantically equivalent prompts contain minor variations. This underscores the importance of improving the robustness of LLMs to natural prompt perturbations.
@@
Existing studies have proposed various strategies to improve the robustness of LLMs, varying from Input-level interventions to model-augmentation approaches. Input-level interventions (Wang et al., 2024a; Agrawal et al., 2025) aim to sanitize or rephrase perturbed prompts before passing them into the model. For example, LLMs can be used to denoise inputs or generate multiple paraphrased variants, from which the most effective is selected. Although these techniques improve performance under prompt perturbations, they do not modify the internal robustness of the model itself. In contrast, model-augmentation approaches (Hu et al., 2024; Agrawal et al., 2025; Sun et al., 2023) add trainable components such as soft prompts, retrieval-augmented generation (RAG), or Low-Rank Adaptation (LoRA) to handle input variability. However, these methods increase system complexity and require additional training efforts. These limitations underscore the need for a lightweight method to directly enhance the inherent robustness of LLMs without extensive retraining or architectural modifications.
@@
In recent years, knowledge editing techniques have been proposed for LLMs (Li et al., 2024d; Wang et al., 2024c; Zhang et al., 2024), enabling efficient post-training updates without full model retraining. Therefore, it is an intuitive idea to explore whether such localized modifications can also improve the robustness of CodeLLMs. However, existing knowledge editing methods mainly tackle factual knowledge (Meng et al., 2022a; Li et al., 2024c; Dai et al., 2021; Wu et al., 2023). They depend on the subject tokens or specific phrases in a single sentence to locate the areas for editing. In contrast, robustness-oriented tasks often involve complex, multi-sentence natural language prompts, making it more difficult to identify meaningful intervention targets. DINM (Wang et al., 2024b) recently applies knowledge editing to detoxification tasks. While the inputs in these tasks are also complex, the availability of gold-standard safe responses enables direct supervision during editing. In contrast, code generation tasks lack a single correct output, presenting unique challenges for applying knowledge editing in this context.
@@
To assess the effectiveness of the proposed framework, we conduct experiments on two widely used code generation benchmarks: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), along with their perturbed counterparts provided by NLPerturbator (Chen et al., 2024b). These perturbations are designed based on empirical observations of real-world user interactions with code LLMs. We evaluate our method on two representative open-source LLMs: CodeLlama-7b and Qwen2.5-Coder-7B. To provide a comprehensive evaluation, we compare CREME with four strong baselines. These include two robustness-enhancement methods (i.e., Self-Denoising (Agrawal et al., 2025) and LoRA Fine-Tuning (Hu et al., 2024)) and two knowledge-editing approaches (i.e., ROME (Meng et al., 2022a) and DINM (Wang et al., 2024b)). Experimental results demonstrate that ❶ CREME significantly improves model robustness, yielding a 63% relative increase in Pass@1 accuracy on perturbed prompts. ❷ CREME exhibits strong generalization across diverse perturbation types. Editing the model based on a single perturbed instance restores up to 30% of the overall code generation accuracy within that perturbation category. ❸ Causal tracing-based layer localization plays a critical role in robustness enhancement by accurately identifying the robustness-sensitive regions within the model. ❹ Robustness-sensitive layers exhibit a clustering pattern, and their positions shift depending on the model architecture. ❺ CREME maintains stable performance on clean inputs, with accuracy deviations within $\pm$1%.
@@
  • We introduce CREME, a novel editing framework that uses a single pair of original and perturbed prompts to identify robustness-sensitive layers and update targeted parameters to enhance the robustness of LLMs. The replication package of our work is publicly available at (cre, 2025).
@@ 3.2. Key Layer Localization
Setup. We first compute the baseline performance of $P_{ori}$ and $P_{pert}$ using the pass@k metric (Chen et al., 2021) by generating outputs from the model G and checking functional correctness (e.g., via unit tests). Let $\text{Acc}_{ori}$ and $\text{Acc}_{pert}$ denote the pass@1 for $P_{ori}$ and $P_{pert}$.
@@
This ratio captures how much of the accuracy gap between $P_{ori}$ and $P_{pert}$ is recovered by intervention at layer $L_{\ell}$, normalized to the maximum possible improvement. We define the key layer $L_{\ell^{*}}$ as the one with the highest restoration improvement. If multiple layers achieve the highest restoration improvement, we adopt the approach proposed by Wang et al. (Wang et al., 2024b). We compute the Euclidean distance between the hidden states of each candidate layer when processing $P_{ori}$ and $P_{pert}$, and select the layer with the greatest distance.
@@ 3.3. Layer Editing
Following the localization of the most critical layer for the degradation of robustness, we proceed to perform parameter-level editing to enhance the robustness of the model to a specific perturbation type. Each layer’s MLP block in Equation 3 is a two-layer neural network (Dai et al., 2021; Meng et al., 2022a) and the second neural network can be expressed as:
@@
$h_{\ell}^{down}$ denotes the intermediate activation obtained by applying the first linear transformation and nonlinear activation (e.g., GELU) to the input $h_{\ell-1}$, while $W_{\ell}^{V}$ represents the output projection matrix that maps the high-dimensional feedforward features back to the hidden size of the model. Previous work (Dai et al., 2021; Meng et al., 2022a; Wang et al., 2024b; Geva et al., 2022; Lee et al., 2024) has demonstrated that $W_{\ell}^{V}$ plays a critical role in knowledge routing and is often the most effective target for localized interventions. Consequently, we edit $W_{\ell^{*}}^{V}$ in the key layer $L_{\ell^{*}}$ to improve the robustness of the model $G$.
@@ 4. Experimental Setup
[…]
@@ 4.1. Baseline
4.1.1. Self-Denoising (Agrawal et al., 2025)
is a prompt-based technique designed to improve the resilience of LLMs to instruction-level perturbations. The method guides the model itself to recover the clean version of a perturbed input before execution. In our experiments, we employ the iterative variant (SDi), where the model is employed multiple times to progressively reduce noise in the input prompt. We select SDi because it outperforms the standard self-denoising approach (Agrawal et al., 2025). However, it does not improve the inherent robustness of the model but instead relies on optimizing the input prompt. In real-world scenarios, we cannot assume that users will modify or refine their prompts prior to interacting with an LLM.
4.1.2. LoRA Fine-tuning (Agrawal et al., 2025)
@@ 4.2. Dataset
  • HumanEval (Chen et al., 2021) is a benchmark consisting of 164 hand-written Python programming problems. Each problem includes an input prompt, a reference solution, and a set of test cases. The prompt comprises a function signature, a functional description, and several example outputs.
  • MBPP (Austin et al., 2021) consists of 974 Python coding problems, including a manually verified subset of 427 problems curated by the authors. Each problem includes a prompt containing a functional description, a reference solution, and test cases. We leverage the manually verified subset to conduct our study.
Based on a literature review and an online survey of practitioners, Chen et al.’s work (Chen et al., 2024b) categorizes common prompt perturbations and propose an automated framework, NLPerturbator, which applies each type of perturbation to a given set of prompts. We follow their work and get the perturbed dataset (i.e., HumanEval-R and MBPP-R). As shown in Table 4, we adopt 20 perturbation types that span lexical and syntactic variations. These perturbations are designed to preserve the original semantics while introducing surface-level deviations that challenge the robustness of models.
@@ 4.3. Evaluation Metrics
Pass@1 is a standard metric for evaluating the functional correctness of generated code (Chen et al., 2021). It measures the proportion of generated outputs that pass all test cases on the first attempt. In this study, we adopt pass@1 rather than higher-$k$ variants (e.g., pass@5 or pass@10), as it reflects the practical setting in which users expect a correct solution from a single generation without manual selection or reranking. Formally, let $n$ denote the number of generated samples for a given prompt, $c$ the number of correct generations. The unbiased pass@k metric is defined as:
@@ 5. Results and Analysis
[…]
@@
[…]
@@ 5.1. RQ1: Effectiveness
Σ Pass@1 Comparison: CREME exhibits strong potential in enhancing model robustness. As shown in Table 2, it achieves the highest Pass@1 scores in 7 out of 8 settings. In all cases, CREME outperforms the unedited model on perturbed prompts, improving the ability of models to generate functionally correct code despite input perturbations. Notably, the Self-Denoising approach performs poorly in most settings, except for QWenCoder+MBPP, contrasting with its previously reported effectiveness (Agrawal et al., 2025). This discrepancy may be attributed to two factors: ❶ unlike like QwenCoder, CodeLlama is not instruction-tuned, which limits its ability to revise perturbed prompts based on meta-instructions; ❷ prompts in HumanEval average 231.6 characters in length, significantly longer than those in MBPP (88.2 characters), which increases the likelihood of semantic drift during denoising.
@@ 5.2. RQ2: Generality
[…]
@@ 6.2. Costs of CREME
[…]
@@ 7. Related Work
In this section, we present a comprehensive review of prior research on LLM robustness and the knowledge editing methods.
@@ 7.1. LLM Robustness
Despite achieving impressive performance in increasingly sophisticated tasks (Liu et al., 2023; Feng and Chen, 2024; Geng et al., 2024), LLMs remain sensitive to input perturbations. While humans are generally robust to minor variations in natural language task descriptions (Walkington et al., 2019), LLMs often produce significantly different outputs in response to such changes (Moradi and Samwald, 2021; Honarvar et al., 2025), highlighting their limited robustness. Wang et al. (Wang et al., 2021) and Zhu et al. (Zhu et al., 2023) introduce benchmark suites such as Adversarial GLUE and PromptRobust, which systematically evaluate LLMs under adversarial or subtly perturbed prompts.
@@
In the domain of code generation, a series of studies have revealed that current code LLMs are sensitive to even minor variations (Wang et al., 2022; Shirafuji et al., 2023; Lam et al., 2025). Chen et al. (Chen et al., 2024b) categorize 18 types of natural language perturbations along with three co-occurring combinations and develop NLPerturbator, a framework targeting real-world prompt perturbations. Their findings show that such perturbations can substantially degrade code generation performance(e.g., up to 21.2%, and 4.8% to 6.1% on average). Lin et al. (Lin et al., 2025) focus on non-functional robustness and introduce RobuNFR, a benchmark designed to evaluate LLMs under real-world noisy contexts in question-answering (QA) tasks. Their results demonstrate that LLMs are highly sensitive to such noise, thereby extending robustness analysis beyond prompt-level perturbations. Mastropaolo et al. (Mastropaolo et al., 2023) conduct an empirical study on GitHub Copilot, showing that 46% of Copilot’s outputs changed when given semantically equivalent paraphrases. This indicates a concerning lack of robustness in popular commercial models.
@@
To improve robustness, researchers have proposed a variety of mitigation strategies. Agrawal et al.(Agrawal et al., 2025) introduce selective instruction augmentation and inference techniques, including self-denoising and representation alignment. They find that self-denoising achieves substantially higher performance gains than alternative strategies. Hu et al.(Hu et al., 2024) propose two mitigation strategies for retrieval-augmented generation: robust prompt alignment, which maps perturbed prompts to semantically equivalent canonical forms to stabilize retrieval, and retrieval consistency filtering, which filters out prompts yielding inconsistent retrieval results across paraphrases. Wang et al. (Wang et al., 2024a) explore instruction-level noise in instruction-tuned models and find that adversarial prompts can severely degrade task performance. They further propose robustness training using paraphrased instruction variants to improve model resilience. Different from prior studies, we apply knowledge editing to enhance model robustness, without relying on additional ensemble strategies. Our method requires only a single example for intervention, offering a novel approach to improving the robustness of LLMs.

### v3

@@ 1. Introduction
In recent years, the rapid development of large language models (LLMs) has led to the emergence of powerful models, including ChatGPT (achiam2023gpt), LLaMA (touvron2023llama), and DeepSeek (liu2024deepseek). Trained on large-scale textual corpora, these models exhibit strong generalization capabilities and have achieved notable success across a wide range of software engineering tasks. Among these tasks, code generation has emerged as a key application in AI-assisted software engineering, attracting growing research attention (chen2021evaluating; li2023starcoder; touvron2023llama; lin2024llm; chen2024code). Typically, practitioners employ LLMs by providing natural language descriptions and LLMs generate the source code, which automates programming tasks and accelerates development workflows.
@@
The natural language description in a prompt is crucial for conveying the requirements defined by users to LLMs. Prior studies (du2023classeval; chen2021evaluating; austin2021program) have evaluated the code generation capabilities of LLMs using datasets consisting of human-verified prompts. However, in real-world scenarios, prompts submitted to LLMs often differ in wording, syntax, and formatting. They may also contain typographical errors or redundant expressions. Prior studies (shen2023chatgpt; wang2023robustness; chen2024nlperturbator; shirafuji2023exploring) have shown that LLMs are sensitive to such minor variations, even a slight change may lead to a completely different result. Not all users of LLMs are skilled prompt engineers capable of making precise, error-free prompts. Therefore, it is essential to ensure output stability when semantically equivalent prompts contain minor variations. This underscores the importance of improving the robustness of LLMs to natural prompt perturbations.
@@
Existing studies have proposed various strategies to improve the robustness of LLMs, varying from Input-level interventions to model-augmentation approaches. Input-level interventions (wang2024resilience; agrawal2025enhancing) aim to sanitize or rephrase perturbed prompts before passing them into the model. For example, LLMs can be used to denoise inputs or generate multiple paraphrased variants, from which the most effective is selected. Although these techniques improve performance under prompt perturbations, they do not modify the internal robustness of the model itself. In contrast, model-augmentation approaches (hu2024prompt; agrawal2025enhancing; sun2023evaluating) add trainable components such as soft prompts, retrieval-augmented generation (RAG), or Low-Rank Adaptation (LoRA) to handle input variability. However, these methods increase system complexity and require additional training efforts. These limitations underscore the need for a lightweight method to directly enhance the inherent robustness of LLMs without extensive retraining or architectural modifications.
@@
In recent years, knowledge editing techniques have been proposed for LLMs (li2024model; wang2024knowledge; zhang2024comprehensive), enabling efficient post-training updates without full model retraining. Therefore, it is an intuitive idea to explore whether such localized modifications can also improve the robustness of CodeLLMs. However, existing knowledge editing methods mainly tackle factual knowledge (meng2022locating; li2024pmet; dai2021knowledge; wu2023depn). They depend on the subject tokens or specific phrases in a single sentence to locate the areas for editing. In contrast, robustness-oriented tasks often involve complex, multi-sentence natural language prompts, making it more difficult to identify meaningful intervention targets. DINM (wang2024detoxifying) recently applies knowledge editing to detoxification tasks. While the inputs in these tasks are also complex, the availability of gold-standard safe responses enables direct supervision during editing. In contrast, code generation tasks lack a single correct output, presenting unique challenges for applying knowledge editing in this context.
@@
To assess the effectiveness of the proposed framework, we conduct experiments on two widely used code generation benchmarks: HumanEval (chen2021evaluating) and MBPP (austin2021program), along with their perturbed counterparts provided by NLPerturbator (chen2024nlperturbator). These perturbations are designed based on empirical observations of real-world user interactions with code LLMs. We evaluate our method on two representative open-source LLMs: CodeLlama-7b and Qwen2.5-Coder-7B. To provide a comprehensive evaluation, we compare CREME with four strong baselines. These include two robustness-enhancement methods (i.e., Self-Denoising (agrawal2025enhancing) and LoRA Fine-Tuning (hu2024prompt)) and two knowledge-editing approaches (i.e., ROME (meng2022locating) and DINM (wang2024detoxifying)). Experimental results demonstrate that ❶ CREME significantly improves model robustness, yielding a 63% relative increase in Pass@1 accuracy on perturbed prompts. ❷ CREME exhibits strong generalization across diverse perturbation types. Editing the model based on a single perturbed instance restores up to 30% of the overall code generation accuracy within that perturbation category. ❸ Causal tracing-based layer localization plays a critical role in robustness enhancement by accurately identifying the robustness-sensitive regions within the model. ❹ Robustness-sensitive layers exhibit a clustering pattern, and their positions shift depending on the model architecture. ❺ CREME maintains stable performance on clean inputs, with accuracy deviations within $\pm$1%.
@@
  • We introduce CREME, a novel editing framework that uses a single pair of original and perturbed prompts to identify robustness-sensitive layers and update targeted parameters to enhance the robustness of LLMs. The replication package of our work is publicly available at (creme2025).
@@ 3.2. Key Layer Localization
Setup. We first compute the baseline performance of $P_{ori}$ and $P_{pert}$ using the pass@k metric (chen2021evaluating) by generating outputs from the model G and checking functional correctness (e.g., via unit tests). Let $\text{Acc}_{ori}$ and $\text{Acc}_{pert}$ denote the pass@1 for $P_{ori}$ and $P_{pert}$.
@@
This ratio captures how much of the accuracy gap between $P_{ori}$ and $P_{pert}$ is recovered by intervention at layer $L_{\ell}$, normalized to the maximum possible improvement. We define the key layer $L_{\ell^{*}}$ as the one with the highest restoration improvement. If multiple layers achieve the highest restoration improvement, we adopt the approach proposed by Wang et al. (wang2024detoxifying). We compute the Euclidean distance between the hidden states of each candidate layer when processing $P_{ori}$ and $P_{pert}$, and select the layer with the greatest distance.
@@ 3.3. Layer Editing
Following the localization of the most critical layer for the degradation of robustness, we proceed to perform parameter-level editing to enhance the robustness of the model to a specific perturbation type. Each layer’s MLP block in Equation 3 is a two-layer neural network (dai2021knowledge; meng2022locating) and the second neural network can be expressed as:
@@
$h_{\ell}^{down}$ denotes the intermediate activation obtained by applying the first linear transformation and nonlinear activation (e.g., GELU) to the input $h_{\ell-1}$, while $W_{\ell}^{V}$ represents the output projection matrix that maps the high-dimensional feedforward features back to the hidden size of the model. Previous work (dai2021knowledge; meng2022locating; wang2024detoxifying; geva2022transformer; lee2024mechanistic) has demonstrated that $W_{\ell}^{V}$ plays a critical role in knowledge routing and is often the most effective target for localized interventions. Consequently, we edit $W_{\ell^{*}}^{V}$ in the key layer $L_{\ell^{*}}$ to improve the robustness of the model $G$.
@@ 4. Experimental Setup
Addition (A1 - A3)
A1 Extra Space inside Words $\rightarrow Write a python function to ch eck$
A2 Repeated Words Write a python function to check … $\rightarrow$ Write a python python function to check …
A3 Repeated Chars Write a python function to check … $\rightarrow$ Write a pyython function to check …
Deletion (D1 - D4)
D1 Char Deletion Write a python function to check … $\rightarrow$ Write a pthon function to check …
D2 Preposition Deletion Write a python function to check … $\rightarrow$ Write a python function_check …
D3 Determiner Deletion … to check if the given number is … $\rightarrow$ … to check if_given number is …
D4 Space Deletion Write a python function to check … $\rightarrow$ Write a python functionto check …
Editing (E1 - E6)
@@
E3 Grammatical Person Variation Write a python function to check … $\rightarrow$ Writes a python function to check …
E4 Active/Passive Voice Variation Write a python function to check … $\rightarrow$ A python function is written to check …
E5 Word Class Variation Write a python function to check … $\rightarrow$ Writer a python function to check …
@@
Swap (S1 - S2)
S1 Swap Adjacent Chars Write a python function to check … $\rightarrow$ Write a python fucntion to check …
S2 Swap Adjacent Words Write a python function to check … $\rightarrow$ Write a python to function check …
Paraphrasing (P1 - P2)
@@
P2 Declarative to Interrogative Print even numbers from a list of numbers. $\rightarrow$ Can you print even numbers from a list of numbers?
Combinations (C1 - C3)
C1 A1 + E1 Write a python function to check … $\rightarrow$ Write a py thon fanction to check …
@@ 4.1. Baseline
4.1.1. Self-Denoising (agrawal2025enhancing)
is a prompt-based technique designed to improve the resilience of LLMs to instruction-level perturbations. The method guides the model itself to recover the clean version of a perturbed input before execution. In our experiments, we employ the iterative variant (SDi), where the model is employed multiple times to progressively reduce noise in the input prompt. We select SDi because it outperforms the standard self-denoising approach (agrawal2025enhancing). However, it does not improve the inherent robustness of the model but instead relies on optimizing the input prompt. In real-world scenarios, we cannot assume that users will modify or refine their prompts prior to interacting with an LLM.
4.1.2. LoRA Fine-tuning (agrawal2025enhancing)
@@ 4.2. Dataset
  • HumanEval (chen2021evaluating) is a benchmark consisting of 164 hand-written Python programming problems. Each problem includes an input prompt, a reference solution, and a set of test cases. The prompt comprises a function signature, a functional description, and several example outputs.
  • MBPP (austin2021program) consists of 974 Python coding problems, including a manually verified subset of 427 problems curated by the authors. Each problem includes a prompt containing a functional description, a reference solution, and test cases. We leverage the manually verified subset to conduct our study.
Based on a literature review and an online survey of practitioners, Chen et al.’s work (chen2024nlperturbator) categorizes common prompt perturbations and propose an automated framework, NLPerturbator, which applies each type of perturbation to a given set of prompts. We follow their work and get the perturbed dataset (i.e., HumanEval-R and MBPP-R). As shown in Table 1, we adopt 20 perturbation types that span lexical and syntactic variations. These perturbations are designed to preserve the original semantics while introducing surface-level deviations that challenge the robustness of models.
@@ 4.3. Evaluation Metrics
Pass@1 is a standard metric for evaluating the functional correctness of generated code (chen2021evaluating). It measures the proportion of generated outputs that pass all test cases on the first attempt. In this study, we adopt pass@1 rather than higher-$k$ variants (e.g., pass@5 or pass@10), as it reflects the practical setting in which users expect a correct solution from a single generation without manual selection or reranking. Formally, let $n$ denote the number of generated samples for a given prompt, $c$ the number of correct generations. The unbiased pass@k metric is defined as:
@@ 5. Results and Analysis
CodeLlama + HumanEval CodeLlama + MBPP QWenCoder + HumanEval QWenCoder + MBPP
Method C3↑ D2↑ C3↑ D2↑ C3↑ D2↑ C3 D2↑ Avg↑
Base LLM (non-perturbation) 0.600 0.583 0.688 0.710 0.739 0.625 0.780 0.769 -
Base LLM (perturbation) 0.220 0.150 0.208 0.190 0.300 0.238 0.190 0.375 -
@@
CodeLlama + HumanEval CodeLlama + MBPP QWenCoder + HumanEval QWenCoder + MBPP
Method C3↑ D2↑ C3↑ D2↑ C3↑ D2↑ C3 D2↑ Avg↑
@@ 5.1. RQ1: Effectiveness
Σ Pass@1 Comparison: CREME exhibits strong potential in enhancing model robustness. As shown in Table 2, it achieves the highest Pass@1 scores in 7 out of 8 settings. In all cases, CREME outperforms the unedited model on perturbed prompts, improving the ability of models to generate functionally correct code despite input perturbations. Notably, the Self-Denoising approach performs poorly in most settings, except for QWenCoder+MBPP, contrasting with its previously reported effectiveness (agrawal2025enhancing). This discrepancy may be attributed to two factors: ❶ unlike like QwenCoder, CodeLlama is not instruction-tuned, which limits its ability to revise perturbed prompts based on meta-instructions; ❷ prompts in HumanEval average 231.6 characters in length, significantly longer than those in MBPP (88.2 characters), which increases the likelihood of semantic drift during denoising.
@@ 5.2. RQ2: Generality
CodeLlama + HumanEval CodeLlama + MBPP QWenCoder + HumanEval QWenCoder + MBPP
Method C3 D2 C3 D2 C3 D2 C3 D2
@@ 6.2. Costs of CREME
Method Localization Time (s) Train Time (s) Test Time (s)
@@ 7. Related Work
In this section, we present a comprehensive review of prior research on LLM robustness and the knowledge editing methods. Beyond LLM robustness, SE work has studied software security and practical deployment issues (liu2025empirical; li2022poison; chen2025diffploitfacilitatingcrossversionexploit; pan2022automated; yang2024federated).
@@ 7.1. LLM Robustness
Despite achieving impressive performance in increasingly sophisticated tasks (liu2023fill; feng2024prompting; geng2024large), LLMs remain sensitive to input perturbations. While humans are generally robust to minor variations in natural language task descriptions (walkington2019effect), LLMs often produce significantly different outputs in response to such changes (moradi2021evaluating; honarvar2025turbulence), highlighting their limited robustness. Wang et al. (wang2021adversarial) and Zhu et al. (zhu2023promptrobust) introduce benchmark suites such as Adversarial GLUE and PromptRobust, which systematically evaluate LLMs under adversarial or subtly perturbed prompts.
@@
In the domain of code generation, a series of studies have revealed that current code LLMs are sensitive to even minor variations (wang2022recode; shirafuji2023exploring; lam2025codecrash). Chen et al. (chen2024nlperturbator) categorize 18 types of natural language perturbations along with three co-occurring combinations and develop NLPerturbator, a framework targeting real-world prompt perturbations. Their findings show that such perturbations can substantially degrade code generation performance(e.g., up to 21.2%, and 4.8% to 6.1% on average). Lin et al. (lin2025robunfr) focus on non-functional robustness and introduce RobuNFR, a benchmark designed to evaluate LLMs under real-world noisy contexts in question-answering (QA) tasks. Their results demonstrate that LLMs are highly sensitive to such noise, thereby extending robustness analysis beyond prompt-level perturbations. Mastropaolo et al. (mastropaolo2023robustness) conduct an empirical study on GitHub Copilot, showing that 46% of Copilot’s outputs changed when given semantically equivalent paraphrases. This indicates a concerning lack of robustness in popular commercial models.
@@
To improve robustness, researchers have proposed a variety of mitigation strategies. Agrawal et al.(agrawal2025enhancing) introduce selective instruction augmentation and inference techniques, including self-denoising and representation alignment. They find that self-denoising achieves substantially higher performance gains than alternative strategies. Hu et al.(hu2024prompt) propose two mitigation strategies for retrieval-augmented generation: robust prompt alignment, which maps perturbed prompts to semantically equivalent canonical forms to stabilize retrieval, and retrieval consistency filtering, which filters out prompts yielding inconsistent retrieval results across paraphrases. Wang et al. (wang2024resilience) explore instruction-level noise in instruction-tuned models and find that adversarial prompts can severely degrade task performance. They further propose robustness training using paraphrased instruction variants to improve model resilience. Different from prior studies, we apply knowledge editing to enhance model robustness, without relying on additional ensemble strategies. Our method requires only a single example for intervention, offering a novel approach to improving the robustness of LLMs.